OpenStudy (anonymous):

8TH GRADE MATH HELP!

OpenStudy (anonymous):

How many solutions does the equation have? 2(2X + 5) = 4(X + 3) A. one solution B. infinite solutions C. no solution

OpenStudy (anonymous):

Ok first you need to find what you need to do with the multiplacation

OpenStudy (anonymous):

i mean the perenthesease ;)

OpenStudy (anonymous):

Idk what to do with it..

OpenStudy (anonymous):

you need to multiply the numbers around them

OpenStudy (anonymous):

Like 3 * 4?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

2 x 2

OpenStudy (anonymous):

like 2 *2X

OpenStudy (anonymous):

^

OpenStudy (anonymous):

2 x 2x = 4x

OpenStudy (anonymous):

then 2 * 5

OpenStudy (anonymous):

so you would get what?

OpenStudy (anonymous):

10?

OpenStudy (anonymous):

yes and?

OpenStudy (anonymous):

I have 4x and 10..

OpenStudy (anonymous):

good now do the same for the other 1

OpenStudy (anonymous):

4 * 3 = 12?

OpenStudy (anonymous):

yaaas

OpenStudy (anonymous):

what do you think 4 * X is

OpenStudy (anonymous):

Uhh 3?

OpenStudy (anonymous):

no 4*X

OpenStudy (anonymous):

I don't know..

OpenStudy (anonymous):

It's 4X

OpenStudy (anonymous):

OH

OpenStudy (anonymous):

so you have 4X+10 and 4X+12 what do you think you should do with them

OpenStudy (anonymous):

Add 4x and 4x and 10 and 12?

OpenStudy (anonymous):

no you subtract 4X and 4X and what do you get

OpenStudy (anonymous):

Just X?

OpenStudy (anonymous):

no there is no X see X itself has a 1 in front of it every time and when you have no x you have no solution

OpenStudy (anonymous):

So it would be 1 or nothing?

OpenStudy (anonymous):

It's confusing..

OpenStudy (anonymous):

it has no solution thats the answer

OpenStudy (anonymous):

so its C

OpenStudy (anonymous):

OH I see! I keep forgetting that I'm not suppose to actually be solving it..

OpenStudy (anonymous):

Can you help with another? I have five of these questions

OpenStudy (anonymous):

@tyler.ford

OpenStudy (anonymous):

How many solutions does the equation have? 6 - a = 1/3(9a - 18)

OpenStudy (anonymous):

A. one solution B. infinite solutions C. no solution

OpenStudy (anonymous):

1/3 is a fraction not dividing

OpenStudy (anonymous):

@TheSmartOne

TheSmartOne (thesmartone):

you tag me to a really long post and I have no idea what you need help with

TheSmartOne (thesmartone):

please keep one question per post open up a new post with your question and tag me again :)

OpenStudy (anonymous):

Okay